OpenStudy (anonymous):

How to find the major and minor axis in the equation x^2+4y^2=36

OpenStudy (irishboy123):

\( \large x^2+4y^2=36\) \( \large \frac {x^2}{36} + \frac{4 y^2}{36} = 1\) \( \large (\frac {x}{6})^2 + (\frac{ y}{3})^2 = 1\)